iiMaddii:

The solution to 4.2x = 19.32 is x = ___.

supie:

Just divide both sides by 4.2

MidNiGhtSkY:

do 4.2 times 19.32

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @MidNiGhtSkY do 4.2 times 19.32 \(\color{#0cbb34}{\text{End of Quote}}\) No.

MidNiGhtSkY:

it says 19.32 is x though

nevaehhh:

All you have to do it divide both sides by 4.3 4.2x=19.32



4.2 4.2 x= ???

nevaehhh:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @nevaehhh All you have to do it divide both sides by 4.3 4.2x=19.32 4.2 4.2 x= ??? \(\color{#0cbb34}{\text{End of Quote}}\) those were supposed to be division marks amigos lmaoo-

MidNiGhtSkY:

4.2(19.32) he said 19.32 is x- :>

MidNiGhtSkY:

but ight.

supie:

\(\LARGE\frac{4.2x}{4.2}=\frac{19.32}{4.2}\) \(\cancel{\frac{4.2x}{4.2}}\) cancels all you have left is \(\frac{19.32}{4.2}\) simplify \(\frac{19.32}{4.2}\) then once you get what \(\frac{19.32}{4.2}\) equals then you have your answer \(\frac{19.32}{4.2}=?\) \(x=?\)

MidNiGhtSkY:

you forgot the x-

MidNiGhtSkY:

there is still an x

supie:

Where?

MidNiGhtSkY:

4.2x/4.2 there is still an x-

supie:

There is an x.

MidNiGhtSkY:

x/1 :>

supie:

?

supie:

We were dividing both sides by 4.2

MidNiGhtSkY:

mhm- and so it would be x=19.32/4.2

MidNiGhtSkY:

cause you forgot the x- :>

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @MidNiGhtSkY mhm- and so it would be x=19.32/4.2 \(\color{#0cbb34}{\text{End of Quote}}\) You're saying that is the overall answer?

MidNiGhtSkY:

well- unless it needs to be simplified-

supie:

Oh I see what you're saying we have to divide both sides by 4.2 and there is two sides (both) so 4.2x/4.2 = 19.32/4.2 4.2x/4.2 cancels so all we have left is 19.32/4.2 So that needs to be simplified to get what x equals

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @MidNiGhtSkY well- unless it needs to be simplified- \(\color{#0cbb34}{\text{End of Quote}}\) And yes it needs to be simplified.

MidNiGhtSkY:

i know that- but- nvm-

supie:

No, you can say what you were going to say. I'm listening.

MidNiGhtSkY:

but you never cancelled the x-